# Dedelayed: Deleting remote inference delay via on-device correction

### v1

@@ 5 Experiments
Our experiments utilize the BDD100K video dataset (Yu et al., 2020), containing video of driving scenes at 30 frames per second (fps). Since the dataset does not provide dense segmentation labels for all video frames, we generated pseudo-labels using a pretrained EoMT (Kerssies et al., 2025) model, ignoring pixels with low confidence. We use a subset of 19 labels from Cityscapes.
Our evaluation covers a range of realistic network delays, from 0 to 5 frames, corresponding to 0 to 165 ms. This range is representative of typical round-trip latencies and is sufficient to demonstrate the degradation of conventional remote inference and the resilience of Dedelayed. At training time, the delay $\tau$ was sampled per batch from a uniform distribution over this range. This delay, together with the relevant frames from the video, was fed into the remote model. To replicate real-world usage, we applied compression to the uplink video streams. We chose the resolution and frame rate to fit within reasonable uplink capacity using the lossy WebP image codec at quality 85.
@@
We adopt a multi-stage training strategy, as detailed in Table˜2. The remote and local models are first trained individually and then later combined. Each model is pretrained on the large-scale ImageNet dataset (Russakovsky et al., 2015) for classification, then on the image segmentation task on Cityscapes (Cordts et al., 2016), before being fine-tuned on the smaller BDD100K driving dataset. We specifically train the remote model to have predictive capability by supplying it with a delay-aware (DA) objective: to predict the labels of future frames, conditioned on a tunable delay. By training in stages, we are able to provide guarantees on baseline remote and local model performance. Finally, the remote and local models are glued together, and the full system is jointly fine-tuned on the delay-aware task of streaming semantic segmentation on video.
We train using cross-entropy loss, the Adan (Xie et al., 2024) optimizer, a trapezoidal cosine learning rate schedule, gradient clipping, and selectively applying discriminative fine-tuning or layer-wise learning rate decay (LLRD) (Howard & Ruder, 2018).
Table 2: Training stages for remote, local, and fusion models.
Model Stage Obj. Data Epochs Res. Freeze
Remote (image only)
SegFormer-B5 pre- DU IN1K, CS
SegFormer-B5 1 DU BDD 15 496
@@
MSTransformer2D 1 DU IN1K 320 224
MSTransformer2D 2 DU CS 80 336 bkbn + proj
MSTransformer2D 3 DU CS 80 336 LLRD 0.9
MSTransformer2D 4 DU BDD 15 496 LLRD 0.9
Fusion (local image + remote video-predictive)
MSTransformer2D + EfficientViT-Seg-L1+ViT3D 1 DA BDD 10 480/720 remote img-bkbn + 3D
Stage: pre- = pretrained (external source).
Obj.: DA = delay-aware objective; DU = delay-unaware objective.
@@
Freeze: bkbn + proj = backbone & MS-projections frozen; remote img-bkbn + 3D = remote image backbone and remote 3D encoder frozen; LLRD = layer-wise learning rate decay.
@@ 6 Results
We compare how various inference systems perform under the effect of communication network latency. Fig.˜6 visualizes the various local-only, remote-only, and fused local+remote methods from different stages of our training. Each baseline is effectively an ablation of our final system.
@@
  • Local + remote predictive represents a Dedelayed system, and is thus able to further sustain accuracy by merging the remote predictive features with fresh local features.
@@
Figure 7: Segmentation accuracy (mIoU) versus total latency (milliseconds or frames) for selected local model delays. Points are faded as round-trip latency increases. 4 and 8 ms were interpolated.
While the above results demonstrate the dominance that Dedelayed is capable of, one should also account for the impact of local inference delay. We show this in Fig.˜7 by assessing accuracy versus total latency. For local inference delays of $\leq$8 ms, the “local + remote predictive” method is consistently better across all network round-trip delays in terms of both accuracy and total latency.
@@ 7 Conclusion
Appendix A Appendix
A.1 Code
The fused “local + remote predictive” model is defined below:
@@ A.1 Code
“”“Any 2D image segmentation model. We use one with T1, T2, T3 blocks.”“”
To train or evaluate the models, simply feed in frames separated by the appropriate delay and compute cross-entropy loss or mIoU in the typical fashion.
A.2 Effect of delay jitter
We evaluate how our model performs under delay jitter, i.e., when the delay varies for each frame.
Our model is trained only to maximize accuracy for a fixed, tunable delay input and is not explicitly optimized for jitter. Nonetheless, temporal structure in the data allows it to retain accuracy even when the delay input differs from the observed delay. Fig.˜8 characterizes this, showing performance under different observed delays when the model is force-fed a possibly inaccurate delay as input. Accuracy peaks when the delay input matches the observed delay, and the drop is less sharp when the observed delay exceeds the delay input. The latter might be attributed to selection bias—if the predicted features are relevant for a future frame, they often remain useful for subsequent frames.
Figure 8: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames). Here, the remote model’s delay input has been force-fed a specific value.
In Fig.˜9, we report the above results in matrix form. Although our model was not explicitly trained for mismatched delays or delays beyond 5 frames, it continues to perform well under these conditions. This matrix can also be used to estimate the performance under delay jitter. We model the observed delay as $\tau_{\text{obs}}\sim\mathcal{N}(\mu=\tau_{\text{in}},\,\sigma^{2})$, centered at the delay input to the model, $\tau_{\text{in}}$. The expected accuracy is obtained by taking a weighted sum over mIoU values for each observed delay, using a discretely binned normal probability mass function, with out-of-bounds mass assigned to the boundary bins. The resulting performance is shown in Fig.˜10 for various values of $\sigma$. In many networks, $\sigma=5\text{ ms}$ and $\sigma=15\text{ ms}$ correspond to relatively high jitter. Yet, even for higher values of $\sigma$, our method’s performance maintains its advantage, even assuming optimistically that the competing baselines experience no jitter. This shows that our method performs stably in realistic networks. For comparison, traditional remote inference loses 3.4% mIoU within the first frame of delay, a drop far larger than the degradation our method incurs with the corresponding $\sigma$.
Figure 9: Segmentation accuracy (mIoU) over observed delay and model delay input.
Figure 10: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames). Latency jitter is modeled as a normal distribution.
A.3 Local input resolution
We evaluate on various local input resolutions. Our remote-assisted local model is able to operate at far lower resolutions without losing accuracy.
Figure 11: Segmentation accuracy (mIoU) versus round-trip latency (milliseconds or frames). Further fine-tuned and evaluated on various local input resolutions.

### v2

@@ 6 Evaluation
We evaluate Dedelayed on the task of real-time semantic segmentation of driving scenes using the BDD100K video dataset [dataset_bdd100k], containing video of driving scenes at 30 frames per second (fps). We use the standard 19 label Cityscapes taxonomy [dataset_Cordts2016Cityscapes] for semantic segmentation of urban scenes. Since the BDD100K dataset does not provide dense segmentation labels for all video frames, we generate pseudo-labels using two models. For the validation set, we use the pre-trained DepthAnything Cityscapes segmentation model [yang2024depth_anything_v1]. To our knowledge, this is the most accurate publicly available semantic segmentation model for the Cityscapes taxonomy. For the training set, we use EoMT [kerssies2025eomtvitsecretlyimagesegmentation], which provides high accuracy, but is significantly faster for labeling 70k training images.
We evaluate the performance subject to known delays ranging from 0 to 5 frames, corresponding to 0 to 167 ms at 30 fps. At training time, the delay $\tau$ is sampled per batch from a uniform distribution over this range.
@@
We adopt a multi-stage training strategy, as detailed in Tab.˜2. The remote and local models are first trained individually and then later combined. Each model is pretrained on the large-scale ImageNet dataset [dataset_ILSVRC15] for classification, then on the image segmentation task on Cityscapes [dataset_Cordts2016Cityscapes], before being fine-tuned on the smaller BDD100K driving dataset. We train the remote model to have temporally predictive capability by supplying it with a delay-aware (DA) objective: to predict the labels of future frames conditioned on the degree of delay. During the final training stage, the task-specific layers of the MLP are replaced with an entropy bottleneck and resolution adapter that allow its outputs to be added to the intermediate activation map of the local model. Using this configuration, the local and remote models are trained jointly for the target scenario of local segmentation incorporating delayed predictions from the remote model. The training loss is per-pixel cross-entropy loss. We use the Adan [xie2024adan] optimizer, a warmup-stable-decay learning rate schedule, gradient clipping, and selectively applying discriminative fine-tuning or layer-wise learning rate decay (LLRD) [howard2018ulmfit_discriminative_finetuning_LLRD].
Table 2: Local and remote model components and training setup.
Stage Local layers Remote layers Data Res.
@@ 2 ViT2D, ViT3D BDD 496
3 CNN2D, CoAt2D IN1K 224
4 CNN2D, CoAt2D CS 336
5 CNN2D, CoAt2D BDD 496
Dedelayed
6 CNN2D, CoAt2D ViT2D, ViT3D BDD 480/720
@@ 7 Results
We compare how various inference systems perform under the effect of communication network latency. Tab.˜3 shows the segmentation performance (mIoU) for different configurations (local-only, remote-only, and the proposed local+remote system, Dedelayed). Each configuration serves as an ablation on the final design.
@@
  • Local + remote predictive represents a Dedelayed system, and so is able to further sustain accuracy by merging the remote predictive features with fresh local features.
As illustrated in Fig.˜7, the performance of conventional remote inference is significantly higher (roughly 5%pt. mIoU) than local inference if communication delay is low. However, the accuracy degrades rapidly with increasing network delay, reflecting the detriment of using stale predictions for dense prediction tasks involving motion. Beyond 2 frames (67 ms) of round-trip latency, remote inference becomes worse than local inference (8%pt. mIoU lower at 167 ms). This impact is significantly mitigated by adopting a temporal prediction training objective, which extends the utility of remote inference extending to 5 frames (167 ms). Finally, the proposed configuration (Dedelayed) incorporates temporally predictive features from the remote model as side information to the local model, and almost completely eliminates the drop-off in accuracy due to delay. At the maximum tested delay of 5 frames, Dedelayed provides a 6.7%pt. mIoU improvement over the local baseline—a difference that would typically require using a model roughly 10 times larger [cai2023efficientvit, kerssies2025eomtvitsecretlyimagesegmentation].
@@
Tab.˜3 shows the accuracy of various inference systems under various local and remote subsystem delays. For local inference delays of $\leq$8 ms, the “local + remote predictive” method consistently delivers better accuracy across all network round-trip delays.
Table 3: Accuracy (mIoU) given various subsystem delays (ms).
Local delay Remote delay
0 ms 33 ms 67 ms 100 ms 133 ms 167 ms
Remote (video-predictive)
0.655 0.649 0.640 0.627 0.611 0.596
Local (image only)
0 ms 0.588 0.588 0.588 0.588 0.588 0.588
4 ms 0.585 0.585 0.585 0.585 0.585 0.585
8 ms 0.582 0.582 0.582 0.582 0.582 0.582
33 ms 0.562 0.562 0.562 0.562 0.562 0.562
Dedelayed
0 ms 0.661 0.660 0.655 0.652 0.646 0.641
4 ms 0.656 0.655 0.650 0.647 0.642 0.637
8 ms 0.652 0.650 0.646 0.643 0.637 0.632
33 ms 0.624 0.619 0.616 0.612 0.607 0.603
Sub-frame (4 ms, 8 ms) accuracies were lerped between 0 ms and 33 ms.
Remote jitter
We also investigate the impact of remote delay jitter on system accuracy. LABEL:fig:miou_by_delay_matrix reports accuracy for when the predictive remote model features are conditioned using a different delay than what is later observed. Since the local model is stateless, the only thing important to its accuracy is what delay the received remote features were conditioned on. In general, the most recently received remote features provide the largest boost to local model accuracy. If the remote features are extremely stale, or are missing, the local model may simply ignore them during inference.
@@ 8 Conclusion
# Any 2D image segmentation model. We use one with T1, T2, T3 blocks.
9 Appendix
9.1 Code
The fused “local + remote predictive” model is defined above. To train or evaluate the models, frames are separated by the appropriate delay and optimized for per-pixel cross-entropy loss or evaluated using mIoU.